OpenStudy (mayankdevnani):

How many electrons are present in 1.6 g of methane?

OpenStudy (mayankdevnani):

@amriju @heena @ParthKohli @hartnn @.Sam.

OpenStudy (anonymous):

first do this step n(mole)=mass/molecular weight

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

methane= \[CH_{4}\]

OpenStudy (mayankdevnani):

right @heena

OpenStudy (anonymous):

ya

OpenStudy (mayankdevnani):

so gram molecular mass=12+4=16g

OpenStudy (mayankdevnani):

right

OpenStudy (anonymous):

ya

OpenStudy (mayankdevnani):

but what is the realtion of electrons in mole

OpenStudy (anonymous):

i already given u the formula put value when u ll get n den n=electron/avogadro num

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

nw solve and show me ur ans

OpenStudy (mayankdevnani):

so no. of moles in 1.6g of methane=1.6/16=0.1

OpenStudy (mayankdevnani):

right

OpenStudy (mayankdevnani):

@heena

OpenStudy (mayankdevnani):

@Callisto @UnkleRhaukus @estudier

OpenStudy (anonymous):

do the work den i ll tell whether u are right or not

OpenStudy (mayankdevnani):

no. of moles of methane in 0.1 mole=0.1*6.022*10^23=6.022*10^22

OpenStudy (mayankdevnani):

right

OpenStudy (mayankdevnani):

one molecule of methane has=6+4=10 electrons

OpenStudy (anonymous):

ye u did it ^.^

OpenStudy (mayankdevnani):

it is the power

OpenStudy (anonymous):

yea ne multiply u an wid 4

OpenStudy (mayankdevnani):

@heena one molecule of methane has=6+4=10 electrons i don't understand this line

OpenStudy (anonymous):

see c num.-->6 outermoct electron it is nt in ur coure yet H4 num. --->4

OpenStudy (mayankdevnani):

ok thnx @heena